A farsighted man uses eyeglasses with a refractive power of 3.09 diopters. Wearing the glasses 0.026 m from his eyes, he is able to read books held no closer than 0.270 m from his eyes. He would like a prescription for contact lenses to serve the same purpose. What is the correct contact lens prescription, in diopters?
Answer #1

from lens law

1/f = 1/di + 1/do

3.09 = 1/di + 1/(0.270 - 0.026)

1/di = -1.00836065574

di = -99.17 cm

N = 99.17 cm + 2.60 cm

N = 101.77 cm

1/f = 1/di + 1/do

1/f = -1/1.0177 + 1/0.270

1/f = +2.72 dipoles